When solutions of AgNO3 and NaOH react, the balanced molecular equation is
2AgNO3(aq) + 2NaOH(aq) Ag2O(s) + 2NaNO3(aq) + H2O(l)

How much Ag2O is produced when 0.200 g of AgNO3 and 0.200 g of NaOH react?
Answer #1
calcualate the moles of AgNO3 and moles of NaOH you started with.

Examine the balanced equation, those two react on a one to one mole ratio, so which was the lowest number? that is the most you can produce. So the balanced equation says you get half that (in moles).
answered by: Caitlin
Answer #2

0.136425g
Answer #3

Ans b) 0.136 g

Number of mol of AgNO3 = mass / molar mass

= 0.200 / 169.9 = 0.0012 mol

mol NaOH = 0.200 / 39.997 = 0.005 mol

Number of mol of AgNO3 is less , so it is the limiting reagent here

Number of mol of Ag2O formed = 0.0012 /2 = 0.00059

Mass of Ag2O = mol x molar mass

= 0.00059 x 231.7

= 0.136 grams